OpenStudy (studydino1201):

Solve. 5 = - (-z+3) A. -8 B.8 C.-2 D.2

OpenStudy (will.h):

1st use ditrubtive property so therefor 5= z -3 8= z

OpenStudy (studydino1201):

ok thank you can you help with more @Will.H

OpenStudy (will.h):

Yes tag me with new questions

OpenStudy (studydino1201):

Solve. 5y=135 A. -19 B. -27 C. 27 D. 19

OpenStudy (studydino1201):

@Will.H

OpenStudy (will.h):

Divide 5 on both sides And you'd get y= 27 Please open new questions

OpenStudy (studydino1201):

Solve. -5/8c=20 A. -34 B. -32 C. 32 D. 34

OpenStudy (studydino1201):

Solve. -35=7t A. -5 B. -1/5 C. 28 D. 42

OpenStudy (studydino1201):

@Will.H

OpenStudy (giitter):

Let's solve your equation step-by-step. 5=−(−z+3) Step 1: Simplify both sides of the equation. 5=−(−z+3) 5=z+−3(Distribute) 5=z−3 Step 2: Flip the equation. z−3=5 Step 3: Add 3 to both sides. z−3+3=5+3 z=8 Answer: z=8

OpenStudy (will.h):

Good work glitter and studydino I can't answer those if you don't open a new post and close this one

OpenStudy (studydino1201):

ok

OpenStudy (giitter):

Thank you @Will.H